25 g of a sample contains 50% MgCl2 and 50% MgSO4.The % of Mg in the mixture is.

Grade:

2 Answers

Rahul
4 years ago
Given weight of mixture = 25g
Molecular weight of mixture = 120 +95 =215g
Amount of magnesium in mixture = 24/215 * 25 = 2.8g
% of Mg in mixture = 2.8/25 * 100 = 11.16
 
Harshit
3 years ago
Mwt of Mg*2/Mwt of mixture}*25=Weigt of Mg in mixture
%of Mg in mixture={weight of mg in mixture/given weight of mix}*100=22.63%